# Multi-Layer Transformers Gradient Can be Approximated in Almost Linear Time

### v1

@@ C.5 Bounded entries of matrices
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
where the 1st step is from the definition of $c(X)$, the 2nd step comes from basic linear algebra, the 3rd step is because of Lemma C.16, C.17. ∎
@@ D.1 Gradient of $s(X)$
The results from [25] give the gradient of $c(X)$. By chain rule, the gradient of $s(X)$ is equivalent to the gradient of $c(X)$ from [25], since $c(X)=s(X)-B$ where $B$ is a constant matrix.
Lemma D.1 (Gradient of $s(X)_{i_{0},j_{0}}$, Lemma B.16 in [25]).
@@ D.2 Gradient on $T_{i}(X)$
  • Let $\mathsf{Attn}_{i}$ be defined as Definition 1.2.
@@
By Definition 1.2 and Definition C.9, we have
@@ D.4 Matrix view of gradient on $T_{i}(X)$
$(G_{i}(i_{0},j_{0})\cdot\frac{\mathrm{d}s(X)_{i_{0},j_{0}}}{\mathrm{d}X})(i_{0},*)=\underbrace{G_{i}(i_{0},j_{0})}_{1\times 1}\cdot\underbrace{(B_{6}(X)(i_{0},*)+B_{7}(X)(i_{0},*)+B_{8}(X)(i_{0},*)+B_{2}(X)+B_{4}(X))}_{d\times 1}$
@@ D.5 Matrix view of each term in gradient on $T_{i}(X)$
  • Let $W\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2
@@ D.6 Components of gradient on $T_{i}(X)$
  • Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$ be defined as Definition 1.2.
@@ E.1 Fast computation for $B_{6}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@ E.2 Fast computation for $B_{7}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d},B\in\mathbb{R}^{n\times d}$ be defined in Definition 1.2.
@@ E.3 Fast computation for $B_{8}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@ E.4 Fast computation for $B_{2}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{d\times n},W,W_{V}\in\mathbb{R}^{d\times d},B\in\mathbb{R}^{n\times d}$ be defined in Definition 1.2.
@@ E.5 Fast computation for $B_{4}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@ E.6 Putting everything together
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition 1.2.
@@ F.1 Key concepts
Note that the $A_{1},A_{2}$ matrices in Definition F.1 is $X$ in our setting. Since in [6], they consider a more general setting, where $A_{1},A_{2}$ can be difference matrices, while in our problem, we consider self-attention. Therefore, in our paper, we have $A_{1}=A_{2}=X$.
@@ F.2 Gradient of $s(X)$ on $W$
[6] proved the close form of the gradient of $c(X)=s(X)-B$ with respect to $W$ for a constant matrix $B$. By chain rule, this is equivalent to the gradient of $s(X)$ with respect to $W$.
Lemma F.3 (Gradient of $s(X)$ on $W$, Lemma B.1 in [6]).
@@
  • Let $f(X),h(X),s(X)$ be defined as Definition C.7, C.8, C.9.
  • Let $W\in\mathbb{R}^{d\times d}$ be defined as Definition 1.2. Let $w\in\mathbb{R}^{d^{2}}$ denote the vector representation of $W$.
@@ F.3 Gradient of $L(X)$ on $W$
Differing from the $\ell_{2}$ loss function used in [6], our framework supports arbitrary loss functions. Therefore, we use Lemma F.4 to illustrate the gradient of $L(X)$ on $W$.
@@
  • Let $W\in\mathbb{R}^{d\times d},X\in\mathbb{R}^{n\times d}$ be Defined as Definition 1.2.
@@ F.4 Fast computation
Recall by Lemma C.14, C.15, we have defined $p_{1}(X),p_{2}(X)\in\mathbb{R}^{n\times n}$.
@@ G.2 Gradient of $L(X)$ on $W_{V}$
By Definition C.9 and Definition 1.2, we have
@@ I.1 Tools from previous work
Algorithm 2 Causal attention mask algorithm, Algorithm 4 in [68]
@@ I.2 Fast computation with causal mask
  • Let $\mathsf{Attn}_{i}$ be defined as Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d}$ be defined as Definition 1.2.
@@
  • Let $X\in\mathbb{R}^{n\times d}$ be defined as Definition 1.2.
@@
  • Let $Q,K,V\in\mathbb{R}^{n\times d}$ be defined as Definition 1.2.
@@
  • Let $\mathsf{Attn}(X)$ be defined as Definition 1.2.

### v2

@@ C.5 Bounded entries of matrices
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
where the 1st step is from the definition of $c(X)$, the 2nd step comes from basic linear algebra, the 3rd step is because of Lemma C.17, C.18. ∎
@@ D.1 Gradient of $s(X)$
The results from DSXY [23] give the gradient of $c(X)$. By chain rule, the gradient of $s(X)$ is equivalent to the gradient of $c(X)$ from DSXY [23], since $c(X)=s(X)-B$ where $B$ is a constant matrix.
Lemma D.1 (Gradient of $s(X)_{i_{0},j_{0}}$, Lemma B.16 in DSXY [23]).
@@ D.2 Gradient on $T_{i}(X)$
  • Let $\mathsf{Attn}_{i}$ be defined as Definition C.3.
@@
By Definition C.3 and Definition C.10, we have
@@ D.4 Matrix view of gradient on $T_{i}(X)$
$~{}(G_{i}(i_{0},j_{0})\cdot\frac{\mathrm{d}s(X)_{i_{0},j_{0}}}{\mathrm{d}X})(i_{0},*)$
$=$ $~{}\underbrace{G_{i}(i_{0},j_{0})}_{1\times 1}\cdot\underbrace{(B_{6}(X)(i_{0},*)+B_{7}(X)(i_{0},*)+B_{8}(X)(i_{0},*)+B_{2}(X)+B_{4}(X))}_{d\times 1}$
@@ D.5 Matrix view of each term in gradient on $T_{i}(X)$
  • Let $W\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$ be defined in Definition C.3
@@ D.6 Components of gradient on $T_{i}(X)$
  • Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$ be defined as Definition C.3.
@@ E.1 Fast computation for $B_{6}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@ E.2 Fast computation for $B_{7}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d},B\in\mathbb{R}^{n\times d}$ be defined in Definition C.3.
@@ E.3 Fast computation for $B_{8}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@ E.4 Fast computation for $B_{2}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
  • Let $X\in\mathbb{R}^{d\times n},W,W_{V}\in\mathbb{R}^{d\times d},B\in\mathbb{R}^{n\times d}$ be defined in Definition C.3.
@@ E.5 Fast computation for $B_{4}(X)$ term
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@ E.6 Putting everything together
  • Let $X\in\mathbb{R}^{n\times d},W,W_{V}\in\mathbb{R}^{d\times d}$ be defined in Definition C.3.
@@ F.1 Key concepts
Note that the $A_{1},A_{2}$ matrices in Definition F.1 is $X$ in our setting. Since in AS24a [6], they consider a more general setting, where $A_{1},A_{2}$ can be difference matrices, while in our problem, we consider self-attention. Therefore, in our paper, we have $A_{1}=A_{2}=X$.
@@ F.2 Gradient of $s(X)$ on $W$
AS24a [6] proved the close form of the gradient of $c(X)=s(X)-B$ with respect to $W$ for a constant matrix $B$. By chain rule, this is equivalent to the gradient of $s(X)$ with respect to $W$.
Lemma F.3 (Gradient of $s(X)$ on $W$, Lemma B.1 in AS24a [6]).
@@
  • Let $f(X),h(X),s(X)$ be defined as Definition C.8, C.9, C.10.
  • Let $W\in\mathbb{R}^{d\times d}$ be defined as Definition C.3. Let $w\in\mathbb{R}^{d^{2}}$ denote the vector representation of $W$.
@@ F.3 Gradient of $L(X)$ on $W$
Differing from the $\ell_{2}$ loss function used in AS24a [6], our framework supports arbitrary loss functions. Therefore, we use Lemma F.4 to illustrate the gradient of $L(X)$ on $W$.
@@
  • Let $W\in\mathbb{R}^{d\times d},X\in\mathbb{R}^{n\times d}$ be Defined as Definition C.3.
@@ F.4 Fast computation
Recall by Lemma C.15, C.16, we have defined $p_{1}(X),p_{2}(X)\in\mathbb{R}^{n\times n}$.
@@ G.2 Gradient of $L(X)$ on $W_{V}$
By Definition C.10 and Definition C.3, we have
@@ I.1 Tools from previous work
Algorithm 2 Causal attention mask algorithm, Algorithm 4 in LLS+24d [57]
@@ I.2 Fast computation with causal mask
  • Let $\mathsf{Attn}_{i}$ be defined as Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d}$ be defined as Definition C.3.
@@
  • Let $X\in\mathbb{R}^{n\times d}$ be defined as Definition C.3.
@@
  • Let $Q,K,V\in\mathbb{R}^{n\times d}$ be defined as Definition C.3.
@@
  • Let $\mathsf{Attn}(X)$ be defined as Definition C.3.